# The High W Challenge: Robust Neutrino Energy Estimators for LArTPCs

### v1

@@ IV Response Matrices
the response matrix represents the probability that a neutrino of true energy $E_{\rm true}$ is reconstructed with an estimated energy $E_{\rm est}$. These matrices are normalized such that $\int R(E_{\rm est},E_{\rm true})dE_{\rm est}=1$ regardless of the selection criteria applied to the underlying samples; the total number of events in each matrix thus reflects the degree of inclusiveness required by the corresponding energy estimator. Fig. 2 shows the response matrices for the five energy estimators, using events simulated with GENIE. Response matrices calculated with the other three generators can be found in the Supplemental Material [SupMat].
The matrix for the CCQE-like estimator exhibits a pronounced secondary band offset from the diagonal, primarily arising from multiproton knockout events and final states containing pions—processes neglected by this method. In contrast, the $W^{2}$ estimator, which is based on similar assumptions but explicitly accounts for pion production and multiproton events, shows a single, well-defined diagonal peak.
The proton-based, SF, and calorimetric estimators all feature narrow diagonal peaks and a significant population in the lower-right quadrant. This reflects that these methods essentially reconstruct the time component of the total final-state four-momentum, which equals the true neutrino energy in the absence of FSI, nucleon motion, or undetected particles. The off-diagonal population arises from events affected by FSI, which are challenging to model. The influence of FSI on the different estimators is discussed further in Section V.3. The proton-based and SF estimators also appear more effective at higher energies, as indicated by the narrowing of their distributions above approximately 4 GeV. However, this trend primarily reflects the diminishing number of events to which these estimators can be applied at such energies, implying that highly pure event selections will be required for their effective use in this regime.
@@
(b) $W^{2}$-based estimator.
(c) Proton-based estimator.
(d) SF estimator.
(e) Calorimetric method.
Figure 2: Confusion matrices describing the smearing from true neutrino energy to estimated neutrino energy using each energy estimation.
The fractional error, defined as $(E_{\rm est}-E_{\rm true})/E_{\rm true}$, further illustrates how the choice of estimator and the inclusion of resolution effects influence neutrino energy reconstruction. Figure 3 shows the fractional error distributions for the different estimators obtained using GENIE, assuming perfect kinematic resolutions in Fig. 3(a), and after applying smearing effects in Fig. 3(b). The results for other generators are provided in the Supplementary Material [SupMat]. Assuming perfect kinematic reconstruction, the proton-based, SF, and calorimetric estimators exhibit the best performance. However, the calorimetric method is particularly sensitive to resolution effects associated with energy deposition in LArTPCs, as evident from the comparison between Figs. 3(a) and 3(b). The CCQE-like estimator shows a narrow peak near zero with a pronounced asymmetric tail, corresponding to CCQE and non-CCQE interactions, respectively. A comparison between the CCQE-like and $W^{2}$ estimators highlights the improvement gained by incorporating the invariant mass into the energy reconstruction. Both produce similarly shaped peaks near zero, but the $W^{2}$ method notably lacks the long tail characteristic of the CCQE-like estimator.
@@
(b) After applying resolution effects.
Figure 3: The distribution of fractional error in neutrino energy using the different estimation methods, calculated using events simulated with the GENIE neutrino event generator. All distributions are normalized to 1. The pink dashed line is the calculation assuming overall 20% smearing on the energy calculated with the calorimetric method.
@@ V.1 Estimator Comparisons
Fig. 4 shows the fractional bias and variance obtained using GENIE. Among the estimators, the $W^{2}$ method exhibits the smallest overall bias. This behaviour can be understood by noting that the other estimators yield markedly asymmetric fractional error distributions, as illustrated in Fig. 3. Fig. 4(b) shows the fractional variance in the energy estimators, after including the effects of detector resolution. In the case of the calorimetric method, we show the variance when applying both the individual particle resolutions from Table 2, and the overall 20% smearing (dashed line). This calculation indicates the SF method offers the best variance across most of the energy spectrum. If individual particle resolutions are applied, the calorimetric method performs the best among the more inclusive methods, though its performance suffers considerably if the more conservative 20% overall smearing is applied. This suggests experiments should consider whether to use this method if their calorimetric reconstruction performs well enough. After applying resolution effects, the performance of the SF method, proton based methods offer similar performance, followed by the $W^{2}$ method.
Fig. 4(c) gives the increase in fractional variance for each estimator from applying the resolution effects. This indicates the the SF and proton-based methods are the least susceptible to these effects. The calorimetric method performs similarly, though this again only applies if the most optimistic model of calorimetric reconstruction in the detector is assumed. The $W^{2}$ and CCQE-like estimators are both more susceptible to resolution effects at low energies. In the case of the $W^{2}$ estimator this is because at low energies the hadronic system is often only a single proton and thus the energy estimation is purely driven by the kinematics, whereas the other methods will experience cancellations between the smearing effects to the lepton and hadronic kinematics. In the case of the CCQE-like method the resolution at higher energies is more dominated by the unrealistic assumptions this method makes about the neutrino interaction as opposed to the momentum resolution, and thus the momentum smearing is less impactful.
(a) Fractional bias.
@@
(c) Increase in fractional variance from application of resolution effects.
@@ V.2 Generator Comparisons
Since realistic physics analyses typically employ multiple event generators to evaluate systematic uncertainties, we assess the robustness of each energy estimator with respect to the underlying nuclear model by varying the neutrino event generator and comparing the resulting bias and variance in the reconstructed neutrino energy. Fig. 5 shows the bias as a function of true neutrino energy for each generator. The results for the variance are provided in the Supplementary Material [SupMat].
The benefit of the exclusive selection employed by the SF method emerges here, as the spread between the biases predicted by each generator for this estimator is noticeably smaller than that of the other estimators. It is also interesting to note that the proton-based method does not perform substantially better than the inclusive methods in this test. Of the three inclusive methods, the $W^{2}$ and calorimetric methods offer very similar performance, while the CCQE-like shows a substantial divergence between GiBUU and the other generators above 2 GeV. The predictions from GiBUU disagree with the expectation from the other three generators for all estimators at most energies, expected given GiBUU’s different approach to nuclear transport modeling.
@@
(b) $W^{2}$ method.
(c) Proton based.
(d) SF method.
(e) Calorimetric method.
@@ V.3 Focus on FSI
The CCQE-like method is insensitive to FSI by construction, as it does not rely on information from the hadronic system. However, this conclusion carries the caveat that the factorized treatment separating FSI from the primary interaction is only an approximation of the true nuclear response.
FSI impact both the magnitude and shape of the bias for the $W^{2}$ and calorimetric methods, albeit in different ways. If GENIE is used, FSI has a minimal effect on the $W^{2}$ estimator for neutrino energies above approximately 2 GeV, and the non-FSI bias curve follows a very similar shape to the FSI one for both generators. The proton-based estimator is the most strongly affected estimator, a trend which can be seen more clearly in Figs. 6(c) and 6(d), which show the change in bias when FSI is switched on. The reliance on proton kinematics in this estimator means the visible hadronic energy in the final state is very susceptible to proton-nucleon rescattering inside the nucleus. The change in bias separates the estimators that utilise directional information from those that do not: these shifts are of a very similar shape (and magnitude when calculated with NuWro) for the $W^{2}$ and SF estimators, and likewise for the proton-based and calorimetric estimators. This calculation also exposes a noteworthy difference between the FSI modeling in NuWro and GENIE in that the change in energy bias far larger in NuWro than GENIE for most of the estimation methods.
@@
Figure 6: Fractional bias as a function of true neutrino energy, calculated before and after simulating final state interactions. Dashed lines indicate the calculation without simulating FSI. Panels (b) and (c) show the change in the bias from switching on FSI.
@@
Figure 7: Fractional change in the estimated neutrino energy due to FSI for each estimation method. The CCQE-like method is not displayed.
@@ VI Secondary Variable Dependences
In this section, we use the GENIE generator as our baseline model. For completeness, we repeat the calculations with all other generators, yielding conclusions consistent with GENIE; details are provided in the Supplementary Material [SupMat]. We find that the estimator least sensitive to the choice of generator is the CCQE-like method; however, though this apparent advantage is offset by its consistently large bias and variance across all variables examined.
@@ VI.1 Visible Hadronic Invariant Mass
Fig. 8 shows the bias as a function of the visible hadronic invariant mass  [^1] We do not draw the SF method in this comparison as this is only valid when $W_{\rm vis}=M_{p}$. The first notable feature is the periodic behaviour of the bias observed across all methods, most prominently in the CCQE-like estimator and, to a lesser extent, in the proton-based method. For the CCQE-like, the bias raises and falls as the true hadronic invariant mass approaches values corresponding to integer multiples of the proton mass. This pattern can be interpreted as a larger fraction of the energy being transferred to pions when $W_{\rm vis}$ lies between the peaks in the $W$ spectrum associated with specific proton multiplicities, resulting in greater bias for estimators that neglect pion information. The CCQE-like estimator exhibits a sharp increase in variance once $W>2M_{p}$, marking the boundary beyond which the charged-current quasi-elastic (CCQE) hypothesis underlying this method is no longer valid. All other estimators offer similar performance, with the proton based method fluctuating depending on whether the final state invariant mass is close to an integer number of protons or not. The same periodic behavior in the variance with a period roughly equal to $M_{p}$ is also present, although with a smaller effect.
(a) Fractional bias.
(b) Fractional variance.
@@ VI.2 Missing Hadronic Energy
We define the missing hadronic energy as the sum of the kinetic energies of all neutrons, the kinetic energies of protons and total energies of pions below the detection threshold. Fig. 9 shows the distributions of bias and variance in neutrino energy as a function of missing hadronic energy. The distribution of bias produced by the $W^{2}$ method and CCQE-like methods are flatter than the other estimators, implying greater stability with respect to the variation of missing energy, though this benefit is partially offset by the greater variance shown in Fig. 9(b). It is interesting to note the SF method is still susceptible to the presence of missing energy despite being constructed to infer the its impact through missing transverse momentum. Comparison of this figures with the corresponding bias distributions from the visible invariant mass and lepton angle shows that the missing hadronic energy has the potential to create the largest bias in neutrino energy.
(a) Fractional bias.
(b) Fractional variance.
@@ VI.3 Angular Dependency
One of the features that distinguishes the CCQE-like, $W^{2}$, and SF estimators from the other methods is that they include angular information in their calculation, in contrast with the proton-based and calorimetric methods which purely utilize the kinetic (or total) energies of final state particles, irrespective of their directions. Fig. 10 shows the fractional bias and variance as a function of lepton scattering angle $\theta$. The bias calculation reveals the improvement made to the purely muon kinematics method by accounting for the invariant mass of the full final state, as the muon kinematics increasingly underestimates the neutrino energy as the scattering angle increases. At very large angles the interaction will be highly inelastic making the quasielastic approximation inaccurate, leading to large bias. The $W^{2}$ method overall performs well in this test as the bias is very close to flat as a function of lepton angle, while all other estimators trend downward.
(a) Fractional bias.
(b) Fractional variance.
@@ VII Impact of Estimator Choice on Oscillation Parameters
In this section, we examine how the choice of energy estimator influences oscillation measurements, with the aim of identifying which estimator provides the most robust inference of oscillation parameters. Here, robustness refers to reduced susceptibility to biases arising from the choice of event generator in general and from final-state interactions (FSI) in particular. Since the connection between reconstructed neutrino energy and oscillation parameter inference is non-trivial, the CP-violating phase $\delta_{CP}$ is probed using a toy $\nu_{e}$-appearance analysis at the far detector, while the mass-splitting $\Delta m^{2}_{32}$ is constrained through “shape-only” fits of $\nu_{\mu}$ disappearance spectra. We make no assumption regarding a near detector constraint.

### v2

@@ IV Response Matrices
the response matrix represents the probability that a neutrino of true energy $E_{\rm true}$ is reconstructed with an estimated energy $E_{\rm est}$. These matrices are normalized such that $\int R(E_{\rm est},E_{\rm true})dE_{\rm est}=1$ regardless of the selection criteria applied to the underlying samples; the total number of events in each matrix thus reflects the degree of inclusiveness required by the corresponding energy estimator. Fig. 2 shows the response matrices for the five energy estimators, using events simulated with GENIE. Response matrices calculated with the other three generators can be found in the Supplemental Material [1].
The matrix for the CCQE-like estimator exhibits a pronounced secondary band offset from the diagonal, primarily arising from multiproton knockout events and final states containing pions, processes neglected by this method. We can see the performance of this estimator improves considerably when only applied to events with single protons, shown in Fig. 2(b), or when incorporating the $\Delta$ mass correction in Fig. 2(c). The $W^{2}$ estimator improves on these methods further, with similar performance to the single proton form over a far broader set of final states, and a narrower peak than the $\Delta$ corrected version.
The proton-based, SF, and calorimetric estimators all feature narrow diagonal peaks and a secondary population in the lower-right quadrant. This reflects that these methods essentially reconstruct the time component of the total final-state four-momentum, which equals the true neutrino energy in the absence of final state interactions (FSI), nucleon motion, or undetected particles. The off-diagonal population arises from events affected by FSI, which are challenging to model. The influence of FSI on the different estimators is discussed further in Section V.3.
@@
(b) CCQE-like estimator only applied to single proton events with no pions.
(c) CCQE-like estimator with $\Delta$ mass correction in meson producing events.
(d) $W^{2}$-based estimator.
(e) Proton-based estimator.
(f) SF estimator.
(g) Calorimetric method.
Figure 2: Response matrices describing the smearing from true neutrino energy to estimated neutrino energy using each energy estimation.
The fractional error, defined as $(E_{\rm est}-E_{\rm true})/E_{\rm true}$, further illustrates how the choice of estimator and the inclusion of resolution effects influence neutrino energy reconstruction. Figure 3 shows the fractional error distributions for the different estimators obtained using GENIE, assuming perfect kinematic resolutions in Fig. 3(a), and after applying smearing effects in Fig. 3(c). For the calorimetric method we separately calculate the resolution applying particle-by-particle smearing of the energies used in equation 6 (solid magenta line) and an overall 20% effect (dashed magenta line) to simulate effects impacting the quantity of charge detected by wire planes. The results for other generators are provided in the Supplementary Material [1].
Assuming perfect kinematic reconstruction, the proton-based, SF, and calorimetric estimators exhibit the best performance. After application of resolution effects the best estimator is the SF method, benefiting from the better resolution on the momentum of protons compared with that of pions and the estimators that try to analyze them. The CCQE-like estimator shows a narrow peak near zero with a pronounced asymmetric tail, corresponding to CCQE and non-CCQE interactions, respectively, a feature that is greatly diminished when the estimator is restricted to only single proton events. A comparison between the CCQE-like and $W^{2}$ estimators highlights the improvement gained by incorporating the invariant mass into the energy reconstruction. Both produce similarly shaped peaks near zero, but the $W^{2}$ method notably lacks the long tail characteristic of the CCQE-like estimator.
@@
(b) Shape of fractional bias around zero.
(c) After applying resolution effects.
[…]
@@ V.1 Estimator Comparisons
Fig. 4 shows the fractional bias and variance obtained using GENIE. Among the estimators, the $W^{2}$ method exhibits the smallest overall bias. This behaviour can be understood by noting that the other estimators yield markedly asymmetric fractional error distributions, as illustrated in Fig. 3. Figs. 4(b) and 4(c) shows the fractional variance in the energy estimators, before and after including the effects of detector resolution. In the case of the calorimetric method, we show the variance when applying both the individual particle resolutions from Table 2, and the overall 20% smearing (dashed line). This calculation indicates the SF method offers the best variance across most of the energy spectrum. If individual particle resolutions are applied, the calorimetric method performs the best among the more inclusive methods, though its performance suffers considerably if the more conservative 20% overall smearing is applied. This suggests experiments should consider whether to use this method if their calorimetric reconstruction performs well enough. After applying resolution effects, the performance of the SF method, proton based methods offer similar performance, followed by the $W^{2}$ method.
Fig. 4(d) gives the increase in fractional variance for each estimator from applying the resolution effects. This indicates the the SF and proton-based methods are the least susceptible to these effects. The calorimetric method performs similarly, though this again only applies if the most optimistic model of calorimetric reconstruction in the detector is assumed. The $W^{2}$ and CCQE-like estimators are both more susceptible to resolution effects at low energies. In the case of the $W^{2}$ estimator this is because at low energies the hadronic system is often only a single proton and thus the energy estimation is purely driven by the kinematics, whereas the other methods will experience cancellations between the smearing effects to the lepton and hadronic kinematics. In the case of the CCQE-like method the resolution at higher energies is more dominated by the unrealistic assumptions this method makes about the neutrino interaction as opposed to the momentum resolution, and thus the momentum smearing is less impactful.
(a) Fractional bias. We do not show the full range of bias from the CCQE-like and $\Delta$ corrected CCQE-like methods in order to display the other calculations in more detail.
@@
(c) Fractional variance, including resolution effects.
(d) Increase in fractional variance from application of resolution effects.
@@ V.2 Generator Comparisons
Since realistic physics analyses typically employ multiple event generators to evaluate systematic uncertainties, we assess the robustness of each energy estimator with respect to the underlying nuclear model by varying the neutrino event generator and comparing the resulting bias and variance in the reconstructed neutrino energy. Fig. 5 shows the bias as a function of true neutrino energy for each generator. The results for the variance are provided in the Supplementary Material [1].
The benefits of the exclusive selection employed by the SF method and the modified versions of the CCQE-like estimator emerge here, as the spread between the biases predicted by each generator for these estimators are noticeably smaller than that of the other estimators. It is also interesting to note that the proton-based method does not perform substantially better than the inclusive methods in this test. Of the three inclusive methods, the $W^{2}$ and calorimetric methods offer very similar performance, while the CCQE-like method, with and without the $\Delta$ mass correction, shows a substantial divergence between GiBUU and the other generators above 2 GeV. The predictions from GiBUU disagree with the expectation from the other three generators for most of the estimators across much of the energy range studied, expected given GiBUU’s different approach to nuclear transport modeling.
@@
[…]
(c) CCQE-like with $\Delta$ mass correction.
(d) $W^{2}$ method.
(e) Proton based.
(f) SF method.
(g) Calorimetric method.
@@ V.3 Focus on FSI
The CCQE-like method is insensitive to FSI by construction, as it does not rely on information from the hadronic system. However, this conclusion carries the caveat that the factorized treatment separating FSI from the primary interaction is only an approximation of the true nuclear response. The single proton and $\Delta$ correction variations of the CCQE-like estimator do display some sensitivity to FSI as rescattering of particles can cause events to migrate from one channel to another leading these estimators to examine slightly different populations of events before and after FSI.
FSI impact both the magnitude and shape of the bias for the $W^{2}$ and calorimetric methods, albeit in different ways. If GENIE is used, FSI has a minimal effect on the $W^{2}$ estimator for neutrino energies above approximately 2 GeV, and the non-FSI bias curve follows a very similar shape to the FSI one for both generators. The proton-based estimator is the most strongly affected estimator, a trend which can be seen more clearly in Figs. 6(c) and 6(d), which show the change in bias when FSI is switched on. The reliance on proton kinematics in this estimator means the visible hadronic energy in the final state is very susceptible to proton-nucleon rescattering inside the nucleus. The change in bias separates the estimators that utilise directional information from those that do not: these shifts are of a very similar shape (and magnitude when calculated with NuWro) for the $W^{2}$ and SF estimators, and likewise for the proton-based and calorimetric estimators. This calculation also exposes a noteworthy difference between the FSI modeling in NuWro and GENIE, in that the change in energy bias is far larger in NuWro than GENIE for most of the estimation methods. Figures displaying the variance of each estimator before and after the application of final state interactions can be found in the Supplemental Material [1].
@@
Figure 6: Fractional bias as a function of true neutrino energy, calculated before and after simulating final state interactions. Dashed lines indicate the calculation without simulating FSI. Panels (c) and (d) show the change in the bias from switching on FSI.
@@
The similarity between the proton-based and calorimetric estimators displayed in NuWro is attributed to a lower rate of meson production during FSI compared to that of GENIE. The SF and $W^{2}$ estimators are also very similar, a result of the bias in the $W^{2}$ being a mixture of events with a single proton in the final state, for which the shift in bias has a similar shape to that of the SF method, and a flat contribution from meson producing events. This relationship is studied further in the Supplemental Material [1].
@@
Figure 7: Fractional change in the estimated neutrino energy due to FSI for each estimation method when only considering events surviving selection criteria before and after FSI. The CCQE-like method is not displayed.
@@ VI Secondary Variable Dependences
In this section, we use the GENIE generator as our baseline model. For completeness, we repeat the calculations with all other generators, yielding conclusions consistent with GENIE; details are provided in the Supplementary Material [1]. We find that the estimator least sensitive to the choice of generator is the CCQE-like method; however, though this apparent advantage is offset by its consistently large bias and variance across all variables examined.
@@ VI.1 Visible Hadronic Invariant Mass
Fig. 8 shows the bias as a function of the visible hadronic invariant mass  [^1] We do not draw the SF method or the modified forms of the CCQE-like estimator in this comparison as these is only valid for specific values of $W_{\rm vis}$.. The first notable feature is the periodic behaviour of the bias observed across all methods, most prominently in the CCQE-like estimator and, to a lesser extent, in the proton-based method. For the CCQE-like, the bias raises and falls as the true hadronic invariant mass approaches values corresponding to integer multiples of the proton mass. This pattern can be interpreted as a larger fraction of the energy being transferred to pions when $W_{\rm vis}$ lies between the peaks in the $W$ spectrum associated with specific proton multiplicities, resulting in greater bias for estimators that neglect pion information. The CCQE-like estimator exhibits a sharp increase in variance once $W>2M_{p}$, marking the boundary beyond which the charged-current quasi-elastic (CCQE) hypothesis underlying this method is no longer valid. All other estimators offer similar performance, with the proton based method fluctuating depending on whether the final state invariant mass is close to an integer number of protons or not. The same periodic behavior in the variance with a period roughly equal to $M_{p}$ is also present, although with a smaller effect.
(a) Distribution.
(b) Fractional bias.
(c) Fractional variance.
@@ VI.2 Missing Hadronic Energy
We define the missing hadronic energy as the sum of the kinetic energies of all neutrons (as these are assumed to be invisible), the kinetic energies of all protons below detection threshold, and total energies of pions below the detection threshold. Fig. 9 shows the distributions of bias and variance in neutrino energy as a function of missing hadronic energy. The distribution of bias produced by the $W^{2}$ method and CCQE-like methods are flatter than the other estimators, implying greater stability with respect to the variation of missing energy, though this benefit is partially offset by the greater variance shown in Fig. 9(c). It is interesting to note the SF method is still susceptible to the presence of missing energy despite being constructed to infer the its impact through missing transverse momentum. Comparison of this figure with the corresponding bias distributions from the visible invariant mass and lepton angle shows that the missing hadronic energy has the potential to create the largest bias in neutrino energy.
(a) Distribution.
(b) Fractional bias.
(c) Fractional variance.
@@ VI.3 Angular Dependency
One of the features that distinguishes the CCQE-like, $W^{2}$, and SF estimators from the other methods is that they include angular information in their calculation, in contrast with the proton-based and calorimetric methods which purely utilize the kinetic (or total) energies of final state particles, irrespective of their directions. Fig. 10 shows the fractional bias and variance as a function of lepton scattering angle $\theta$. The bias calculation reveals the improvement made to the purely muon kinematics method by accounting for the invariant mass of the full final state, as the muon kinematics increasingly underestimates the neutrino energy as the scattering angle increases. At very large angles the interaction will be highly inelastic making the quasielastic approximation inaccurate, leading to large bias. We also see a degradation of performance of the estimators according to the variance calculation at large angles. The $W^{2}$ method exhibits a flat bias, while all other estimators trend downward, though this may be offset by the worse resolution than the calorimetric and SF methods displayed by the variance calculation.
(a) Distribution.
(b) Fractional bias.
(c) Fractional variance.
@@
VI.4 Lepton Momentum
Another quantity that is used by several estimators is the energy and/or momentum of the scattered lepton,; a variable that correlates strongly with neutrino energy. In Fig. 11 we study the bias and variance in the estimated neutrino energy as a function of the momentum of the scattered lepton. This shows the $W^{2}$ method has a small, flat bias as a function of this variable, with the next best performing method being the single proton version of the CCQE-like estimator. The SF method offers the lowest variance among the methods compared. The bias and variance both increase for all methods at low muon momentum, as with the lepton scattering angle this is partly due to a larger portion of the energy carried by the neutrino being transferred to the hadronic system in this region of the phase space leading to more complex hadronic processes that are more challenging to simulate and analyze.
(a) Distribution.
(b) Fractional bias.
(c) Fractional variance.
Figure 11: Fractional bias and variance as a function of the scattered muon’s momentum.
@@ VII Impact of Estimator Choice on Oscillation Parameters
In this section, we examine how the choice of energy estimator influences oscillation measurements, with the aim of identifying which estimator provides the most robust inference of oscillation parameters. Here, robustness refers to reduced susceptibility to biases arising from the choice of event generator in general and from final-state interactions (FSI) in particular. Since the connection between reconstructed neutrino energy and oscillation parameter inference is non-trivial, the CP-violating phase $\delta_{CP}$ is probed using a toy $\nu_{e}$-appearance analysis at the far detector, while the mass-splitting $\Delta m^{2}_{23}$ is constrained through “shape-only” fits of $\nu_{\mu}$ disappearance spectra. We make no assumption regarding a near detector constraint.